System of Linear Equations:
Solve the system using the addition method.
2x+3y=-16
5x-10y=30
: System of Linear Equations:
Solve the system using the addition method.
2x+3y=-16
5x-10y=30

Answer by jim_thompson5910(9390)


Start with the given system of equations:
system(2x+3y=-16,5x-10y=30)


5(2x+3y)=5(-16) Multiply the both sides of the first equation by 5.


10x+15y=-80 Distribute and multiply.


-2(5x-10y)=-2(30) Multiply the both sides of the second equation by -2.


-10x+20y=-60 Distribute and multiply.


So we have the new system of equations:
system(10x+15y=-80,-10x+20y=-60)


Now add the equations together. You can do this by simply adding the two left sides and the two right sides separately like this:


(10x+15y)+(-10x+20y)=(-80)+(-60)


(10x+-10x)+(15y+20y)=-80+-60 Group like terms.


0x+35y=-140 Combine like terms. Notice how the x terms cancel out.


35y=-140 Simplify.


y=(-140)/(35) Divide both sides by 35 to isolate y.


y=-4 Reduce.


------------------------------------------------------------------


10x+15y=-80 Now go back to the first equation.


10x+15(-4)=-80 Plug in y=-4.


10x-60=-80 Multiply.


10x=-80+60 Add 60 to both sides.


10x=-20 Combine like terms on the right side.


x=(-20)/(10) Divide both sides by 10 to isolate x.


x=-2 Reduce.


So our answer is x=-2 and y=-4.


Which form the ordered pair .


This means that the system is consistent and independent.


Notice when we graph the equations, we see that they intersect at . So this visually verifies our answer.


drawing(500,500,-12,8,-14,6,<BR>
grid(1),<BR>
graph(500,500,-12,8,-14,6,(-16-2x)/(3),(30-5x)/(-10)),<BR>
circle(-2,-4,0.05),<BR>
circle(-2,-4,0.08),<BR>
circle(-2,-4,0.10)<BR>
) Graph of 2x+3y=-16 (red) and 5x-10y=30 (green)